A random sample of 100 recorded deaths in the United States during the past year showed an average life span of 71.8 years, with a standard deviation of 8.9 years. Does this seem to indicate that the average life span today is greater than 70 years? Use a 0.05 level of significance. The computed value is


1
Expert's answer
2022-05-16T15:18:15-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu\le70"

"H_a:\mu>70"

This corresponds to a right-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a right-tailed test is "z_c = 1.6449."

The rejection region for this right-tailed test is "R = \{z: z> 1.6449\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{71.8-70}{8.9/\sqrt{100}}\approx2.0225"

Since it is observed that "z = 2.0225>1.6449= z_c ," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value is "p=P(Z>2.0225)= 0.021562," and since "p= 0.021562<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is greater than 70, at the "\alpha = 0.05" significance level.